Run Plan for week of Nov15 - Nov22 Updated by Ryan Bayes Nov 15 /------------------------------------------------------------------- Nov 15 (Tuesday) Maintenance day -insert the TEC into beamline /------------------------------------------------------------------- Nov 16 (Wednesday) Maintenance day continues until afternoon After beam returns 1. Characterise the B2 -10G beam -compare directly to run23821 (Be target, B2-10G) - rates - beam spot shape - if run23821 proves insufficient for beam characterisation we may require ~1 hr with the beryllium target to produce a characterisation of a beam comparable to the B2-10G set 2. B2 sweep - use the film degrader to put muon stopping position in the sparce stack. Suggested film degrader setting is position 8 - record the muon stopping position as a function of the - graph as you progress. Take at least 6 data points (more may be helpful) (3. Further suggestions to use the beam through to the morning) /--------------------------------------------------------------------- Nov 17 (Thursday) 1. Quadrupole tests through the day. 2. take "PACT" data. Locating stops in PC 6 High voltage settings to be supplied by Dave /--------------------------------------------------------------------- Nov 18 (Friday) 1. Qudrupole tests as required. 2. Insert 49 hole collimators 3. Start TEC calibration; run with 49 hole collimator. -use spray tune in run21383, that is quads off: SJ=(60,20(-1.8),60(-3.8),100,60(+1.6)),NMR=853.6/926.9 there may be a snap file(?) (continue over the weekend) /--------------------------------------------------------------------- Nov 19 (Saturday) continue TEC calibration runs /--------------------------------------------------------------------- Nov 20 (Sunday) continue TEC calibration runs /--------------------------------------------------------------------- Nov 21 (Monday) end TEC calibration run